The compass of an airplane indicates that it is headed towards north, and its airspeed indicator shows that it is moving through air at 240 km/h. If there is wind of 100 km/h from west to east, what is the velocity of the airplane relative to earth?


1
Expert's answer
2021-07-12T12:11:50-0400
"V=\sqrt{240^2+100^2}=260\frac{km}{h}"